### v1

# How Exclusive are Ethereum Transactions?

Abstract
We analyze 15,097 blocks proposed for inclusion in Ethereum’s blockchain over an 8-minute window on December 3, 2024, during which 38 blocks were added to the chain. We classify transactions as exclusive—present only in blocks from a single builder—or private—absent from the public mempool but included in blocks from multiple builders. We find that exclusive transactions account for 84% of the total fees paid by transactions in winning blocks. Furthermore, we show that exclusivity cannot be fully explained by exclusive relationships between senders and builders: about 7% of all exclusive transactions included on-chain, by value, come from senders who route exclusively to a single builder. Analyzing transaction logs shows that some exclusive transactions are duplicates or variations of the same strategy, but even accounting for that, the share of the total fees paid by transactions in winning blocks is at least 77.2%. Taken together, our findings highlight that exclusive transactions are the dominant source of builder revenues.
1 Introduction
Ethereum separates the roles of the block proposer and the block builder. The proposer is randomly selected from users who have staked at least 32 ETH, and is responsible for adding the next block to the chain. Builders, by contrast, are professional entities that gather transactions from both the public mempool and private channels, and use sophisticated algorithms to decide which transactions to include in a block and in what order. They then compete in an auction by submitting multiple candidate blocks, each with a payment to the proposer in case the block is added to the blockchain. This auction is highly competitive, with thousands of candidate blocks for each block selected for inclusion in Ethereum’s blockchain.
This system, known as Proposer-Builder Separation (PBS), has been effective at selecting the most valuable transactions for inclusion on-chain. Its downside, however, is the extreme concentration in the builders’ market: the top builder (Titan) produces about 50% of blocks, while the top three account for more than 90%. [^1] For up-to-date statistics, see https://www.relayscan.io/. Note that Buildernet is a joint project by Flashbots (who operated their own independent builder until mid-2025) and Beaverbuild (who both participate in Buildernet and run their own builder). This concentration gives a few entities disproportionate influence over which transactions are included, threatening Ethereum’s foundational goals of decentralization and permissionlessness.
The leading explanation for why PBS produced such concentration rests on exclusive transactions and economies of scale: the more often a builder wins, the more attractive it becomes as a destination for order flow that is not shared with competitors, further increasing its wins in future auctions. Quantifying the importance of exclusive transactions, however, is difficult because on-chain data alone cannot reveal whether a transaction was privately shared with multiple builders or only with the eventual winner.
Our paper addresses this challenge using a novel dataset that covers about 28% of all blocks proposed for inclusion in Ethereum during an 8-minute window on December 3, 2024, corresponding to 39 new blocks. [^2] The dataset contains all blocks (winning and non-winning) submitted via the “Agnostic relay” during the study period. They are therefore not a random sample of all proposed blocks (see Section 3 for details and limitations). To our knowledge, together with the companion paper Canidio and Pahari (2025), this is the first study to analyze the content of non-winning Ethereum blocks. It includes 15,097 proposed blocks, 10,793 unique transactions (identified by their hash), and 2,380,014 transaction–block pairings. By comparing the content of proposed blocks with that of the public mempool (sourced from https://mempool.guru/), we classify each transaction as:
  • Public: it appeared in the public mempool;
  • Private: it appeared in blocks from multiple builders but not in the public mempool;
  • Exclusive: it appeared only in blocks from a single builder.
Summary of results.
Of the 5,576 transactions in our dataset that were ultimately included on-chain, 4,124 were public, 810 private, and 642 exclusive. Exclusive transactions, though fewer in number, are by far the most valuable: they account for 84% of all fees paid in winning blocks. This result confirms their central role in block building. We also find that almost all (i.e., 91%) transactions that never make it on-chain are exclusive.
We then study whether transactions change status over time. Among the transactions that first appear as exclusive, 162 remain exclusive during multiple bidding cycles, 174 transition to private in a subsequent bidding cycle, and 12 exclusive transactions later appear in the public mempool. These 12 transactions seem to be user-generated, as they swap on DEX pools using well-known routers (0x and 1inch). Their initial exclusivity is puzzling: ordinary users typically gain no benefit from sending a transaction to a single builder rather than sharing it privately via a known private mempool or broadcasting it publicly. More surprisingly, these transactions later appear in the public mempool, meaning they were first delayed and then exposed to attacks. We speculate that a private mempool operator may have shared them exclusively with one builder, falling back to the public mempool if that builder failed to win. [^3] The presence of exclusive users’ transactions is notable because four of the five major private mempool operators explicitly state that they share transactions with multiple builders (MEV Blocker, Flashbots Protect, Blink, and Merkel; see Janicot and Vinyas, 2025). In contrast, the fifth operator, Metamask Smart Transactions, does not publicly disclose its sharing policies.
We also find cases of transactions changing status within a bidding cycle. We identify 19 cases where transactions start as exclusive to a single builder during the initial part of the auction cycle, but later in the same cycle appear in blocks from other builders. Most of these were submitted via Flashbots Protect: they were initially exclusive to the Flashbots builder before being propagated more broadly.
To investigate the origin of these exclusive transactions, we analyze the relationship between transaction senders and builders. When we simply count transactions, we find that the vast majority of exclusive transactions can be attributed to senders interacting exclusively with one builder. However, when we focus on value and only consider transactions that are included on-chain, according to our preferred operationalization of “transaction sender” only about 7% of value can be attributed to exclusive relationships between searchers and builders, which therefore do not seem very important in explaining the overall importance of exclusive transactions.
We conclude by examining cases where the same sender submits exclusive transactions to multiple builders within a single auction cycle. Using transaction logs, we identify three patterns: (i) 171 transactions have different hashes but are functionally identical to another exclusive transaction in the same cycle, executing the same operation with the same fee; (ii) 494 transactions are identical in execution to another exclusive transaction in the same cycle, but differ in the fee attached; and (iii) 1,179 transactions are similar but not identical to another exclusive transaction in the same cycle, swapping in the same DEX pools but different amounts. While analyzing logs refines our classification of exclusivity, the number of affected transactions is small and does not change our main conclusion.
Relevant literature
To our knowledge, Yang et al. (2024) is the only other paper that examines the content of non-winning blocks. They study builder competition by comparing bids with block revenues (i.e., total payments received by builders), but do not analyze the underlying transactions.
Other strands of the literature explore builders’ competition from different angles. Wu et al. (2024a, b) analyze bidding strategies, both theoretically and via simulation. Titan and Frontier Research (2023); Öz et al. (2024) show that builder profits hinge on how they source transactions, using a classification approach related to ours. Gupta et al. (2023) formalizes how exclusive transactions can drive concentration in the equilibrium of the builders’ market. Finally, Bahrani et al. (2024) model heterogeneous builders and derive conditions under which both builder and proposer markets become concentrated.
Finally, in our companion paper Canidio and Pahari (2025), we use the same dataset to address two related questions: whether block selection matters for users, and competition between arbitrageurs. The paper shows that many user transactions are delayed: although proposed during a bidding cycle, they are not included in the winning block. Furthermore, they also identify two arbitrage bots trading between decentralized (DEX) and centralized (CEX) exchanges. Analyzing the bidding dynamics, they estimate that the implied CEX prices at which they trade USDC/WETH and USDT/WETH are 3.4–4.2 basis points better than the contemporaneous Binance price.
2 Background on PBS
As noted above, in Ethereum, the proposer almost always delegates block construction to builders. Builders collect both user transactions and those submitted by searchers, automated bots that generate transactions in response to market conditions (e.g., price fluctuations) or to other transactions pending in the public mempool. Searchers play a central role in decentralized finance: they engage in arbitrage across DEXs and CEXs, liquidate under-collateralized loans, or exploit temporary price discrepancies. Because their strategies often depend on speed and precise ordering, searchers typically route transactions to builders via private channels. Users can submit transactions in different ways. The default is to broadcast them to the public mempool, where all builders can observe them. This, however, exposes users to attacks. [^4] For example, if a user swaps assets on an Automated Market Maker (AMM), a malicious searcher can mount a “sandwich” attack: front-running the swap with the same trade and then back-running it with the opposite trade. The attacker buys low and sells high, forcing the victim to execute at worse prices. To reduce this risk, many users rely on private mempools, publicly available services which distribute transactions privately to a subset of builders.
The relation between builders and proposers is facilitated by relays. Their primary role is to prevent proposers from altering the submitted blocks. [^5] Technically, this is achieved by requiring proposers to sign an empty block, which is then completed by the relay. For more details, see https://docs.flashbots.net/flashbots-mev-boost/relay. Once a builder creates a block, it submits it to a relay along with a proposed payment to the proposer. The relay then forwards the highest-paying block and its corresponding payment to the proposer upon request. This process resembles an ascending price auction because the relay continuously broadcasts the value of the highest-paying bid—though this broadcast is observed with a delay due to network latency. An auction cycle lasts approximately 12 seconds, after which the winning block is selected, and a new auction begins. During each auction cycle, builders continuously resubmit blocks, modifying the transactions included and bid amounts. Latency plays a crucial role, as proposers are geographically dispersed, and arbitrage transactions often depend on rapid price movements in traditional “off-chain” financial markets. To account for geographic dispersion and fast-moving arbitrage opportunities, multiple relays operate worldwide, with builders submitting to several at once. Our dataset, for instance, comes from a relay called Agnostic Relay.
3 Dataset
Our primary dataset consists of all blocks (winning and non-winning) submitted via Agnostic relay between Ethereum blocks 21,322,622 and 21,322,660 (slots 10,534,387–10,534,425), spanning 2:37:35–2:45:25 PM UTC on December 3, 2024.666Founded in 2022, Agnostic relay held about 20% of the relay market in 2023, declining to around 5% during our study period. We complement this dataset with several public sources:
  • All blocks added to the Ethereum blockchain during our study period, including those not submitted via Agnostic relay. We also track ten additional winning blocks beyond the study period to see whether transactions in our dataset were eventually included on-chain.
  • Hashes of blocks submitted to all major relays: Flashbots, BloXroute, Manifold, Eden, Ultra Sound, SecureRpc, and Aestus. Because any difference in content or bid produces a unique hash, this lets us check whether blocks in our dataset were also sent elsewhere. We find that our 15,097 blocks represent 28.3% of all blocks submitted to major relays during the study period.
  • Data on transactions from the public mempool and from two private mempools operators (MEV Blocker and Flashbots Protect), obtained via https://mempool.guru/ and https://Dune.com.
Finally, we simulate selected transactions in our dataset to test whether transactions with different hashes nonetheless execute the same or similar operations. These simulations are run on an Ethereum Archive Node using the Erigon client, forking at the relevant block heights.
An important detail of PBS for our analysis is that a relay does not run an auction for a given slot if the assigned proposer is not registered with it. This occurs in our data: of the 39 slots we study, 11 show no bidding activity because the proposers were not connected to the Agnostic relay. [^7] These are the slots leading to blocks 21,322,624; 21,322,625; 21,322,627; 21,322,629; 21,322,633; 21,322,634; 21,322,636; 21,322,644; 21,322,646; 21,322,647; and 21,322,651. Also, each block in our dataset has two timestamps: received at (i.e., received from the builder) and made available at (i.e., made available to the proposer). The gap between them reflects the time needed for the relay to simulate the block and verify its validity. [^8] Some blocks are treated “optimistically”—they are made available immediately, with simulation performed in the background. If the simulation later fails, the builder’s next submission is not treated optimistically. Such blocks account for only 3.8% of our sample. This simulation delay is non-trivial: the median is 0.76 seconds, with the 75th percentile at 1.5 seconds. In what follows, we use the received at timestamp, as it better captures the events we study—such as when a builder submits a block or when a transaction first appears.
As noted earlier, our primary dataset covers 15,097 blocks submitted across 28 bidding cycles. The number of blocks per slot ranges from 220 to 951, with a mean of 539 and a median of 509. We can attribute 14,043 of these blocks to 23 known builders. The most active are Titan Builder (7,024 blocks, 46.5%), Rsync Builder (2,259, 15.6%), Flashbots (1,936, 12.8%), and Beaverbuild (1,418, 9.4%). Most blocks in our dataset—12,895 in total—were also submitted to at least one other relay. However, when comparing across all relays, we find an imbalance. Titan Builder submitted 13,679 blocks overall (25.6%), Rsync Builder 7,936 (14.9%), Flashbots 2,626 (4.9%), and Beaverbuild 21,447 (40.2%). Thus, Beaverbuild contributes far fewer blocks to Agnostic Relay than to other relays. We recognize this imbalance as a limitation of our dataset.
Just as searchers and users pay builders to include their transactions, builders pay proposers through a fee reception transaction appended at the end of each block. In our dataset of 15,097 blocks, 13,848 (91.7%) contain such a transaction, corresponding to 12,205 distinct fee reception transactions (different blocks may include the same fee reception transaction). The remaining blocks—those without a fee reception transaction—are low-value submissions in which the builder sets the fee recipient variable directly to the proposer’s address. In these cases, the bid equals the sum of payments sent to that address. In what follows, we use the term “transactions” to mean all transactions excluding fee reception transactions. This yields 10,793 unique transactions (by hash), of which 5,576 were eventually included in a winning block, either in the cycle where they first appeared or later.
4 Results
(a) 11 seconds
(b) 12 seconds
(c) 13 seconds
(d) 14 seconds
Figure 1: Proportion of transactions shared between blocks by different builders, at different moments in the bidding cycle.
As a first step, Figure 1 shows that blocks proposed by different builders contain different transactions. To study these differences, we label transactions using MempoolGuru as either public (it appeared in the public mempool) or non-public (it appeared in a block but not in the mempool). We further split non-public transactions into private (present in blocks from multiple builders) and exclusive (present only in blocks from a single builder).
Focusing on the 5,576 transactions in our dataset that were ultimately included in a winning block, MempoolGuru classifies 4,124 as public and 1,452 as non-public. Of the non-public transactions, 266 were submitted via MEV Blocker, 40 via Flashbots Protect, and 14 through both—indicating that these are private but not exclusive. This leaves 1,132 non-public transactions to classify further. Using our dataset, we find that 642 of them (45% of all non-public transactions in winning blocks) are exclusive: 533 appear only in Titan Builder’s blocks, 107 only in Rsync Builder’s blocks, and 2 only in builder0x69’s blocks.
In value terms, the dominance of exclusives is stark. Of the 14.67 ETH in fees paid by all transactions eventually included in winning blocks, 12.27 ETH (84%) came from exclusive transactions, 1.66 ETH (11%) from private ones, and 0.74 ETH (5%) from public ones. Exclusive transactions are thus by far the builders’ primary source of revenue, reinforcing the idea that they drive the observed dynamics in the builders’ market.
With respect to the 5,217 transactions never included in a winning block, we find 4,769 exclusive transactions and 448 private transactions. Hence, the vast majority of discarded transactions are exclusive.
As noted, a limitation of our analysis is that we observe only a small share of Beaverbuild’s blocks. This creates two concerns. First, some transactions we classify as “exclusive” may in fact also appear in Beaverbuild blocks outside our dataset. For this to occur, a transaction would need to be shared with Beaverbuild and only one other builder in our dataset. However, sharing transactions with exactly two builders is quite rare in our dataset: among the 6,669 non-public transactions in our dataset, we find only 435 that were shared with exactly two builders. Hence, the number of transactions categorized as "exclusive" while in reality they are “private” is probably small. Second, and more importantly, transactions exclusive to Beaverbuild are missing from our analysis whenever Beaverbuild did not submit the corresponding block to Agnostic, even if those transactions were ultimately included on-chain. As a result, the overall share of exclusive transactions during our study window may be larger than what we measure.
4.1 Transactions’ transitions.
We next ask whether transaction status evolves over time. For this purpose, we define each transaction’s status (public, private, or exclusive) within each bidding cycle and then track how that status changes across cycles.
Among the 1,330 transactions observed across multiple bidding cycles, public transactions remain public in all subsequent cycles. Only a handful of private transactions (13 out of 208) later appear in the public mempool. By contrast, the majority of exclusive transactions transition to private or even public. Of the transactions that begin as exclusive, 162 remain exclusive across all observed cycles, while 174 transition to private. [^9] We also find 33 transactions that were initially exclusive to a builder and later included in a subsequent cycle won by the same builder. Because those cycles are not in our primary dataset, we cannot determine whether the transactions remained exclusive or became private. In addition, 12 transactions that were initially exclusive eventually surfaced in the public mempool. Nine of these were present in two cycles and three in three cycles. All appear to be user-generated: three used the 0x router and four the 1inch router. [^10] These transactions are: 0x1daea1584d684385fb25209bad3a49c54a1e2c500e27543334c7b3e695a12ffb, 0x6605c1c36fcd3d696844519d6ff7199fe4ddd34fb4f3e0ed4da71aedfcf1c633, 0x3dadbab8e85b6e48a2c33f9029ea4ebebdd3def8b9d907faff15ef81527de968, 0x9c7364eddb94fe659bc8104585ff86a8a234480fe00a9f9aa3eb410e572ebdf2, 0x9d476dbc524651cbe26bcafab99f6fe40a823a5588fa5cd5002e5b414a8a2f8c, 0x417440b083bdcb529663c6615c32d89a92beb7863f1a5d6b12bca5b94f4a0f87, 0xe56caee27e79dee8a4c7ed6f945ccbc2cc733d761e19bd68329dc364c61e73a7, 0xae5b88bc0bffa04811eda0dfdcd7224599e2757339e18b2d57f4f40d61902a4c, 0xd41139b67a193045b4d60266c084f6dadd0cb95dccd7b2a06ea68653d6a313ec, 0x597fc0cb0610d1e3ea2baf0f6f90d0d741a6429bfdd9b375bfd2493b70d41dcd, 0x44e8105d5c672195877dc1ed970511e0f0a15fc25cbcafe71d51d428d54c215d, 0xd74718e6d8d58bc33334c0c2f6642d9dec49496329329d8fdcb1e698e81abc94. The existence of these transactions is problematic, as they were first delayed and later exposed in the public mempool, leaving users vulnerable to attacks.
A change in transaction status may also occur within a bidding cycle. For instance, a builder unlikely to win may share its exclusive transactions with a competitor that has a better chance. To study this, we focus on Rsync and Titan, the two most prominent builders in our dataset. We define a “new transaction” as one included in a block submitted at least one second after a builder’s initial block, and not present in that first submission. For each such new transaction from Titan and Rsync, we check whether it appeared in another builder’s block at least one second earlier.
We identify 20 cases: 3 involving Rsync and 17 involving Titan. Of Rsync’s three, two were initially exclusive to Beaverbuild and later included in a winning Rsync block. The third—a searcher’s swap—appeared in blocks from five different builders, but only entered Rsync’s block four seconds later (Rsync won that auction).
Of Titan’s 17 cases, 8 came through the Flashbots private mempool. Five originated from a single searcher address1111110x89a99a0a17d37419f99cf8dc5ffa578f3cdb58b5 and appear in Flashbots Protect’s dataset as shared with more than 15 builders. These transactions remained exclusive to Flashbots Builder for about 2.6 seconds before entering Titan’s blocks. Another five transactions from the same address follow the same pattern (initial exclusivity to Flashbots, later appearance in Titan’s blocks) but were never included on-chain. It is plausible that they were also shared via Flashbots Protect, which records only transactions that are eventually included.
In summary, transaction sharing between builders during the bidding cycle does occur but is relatively rare. When it happens, it typically involves transactions sent via Flashbots Protect that are initially exclusive to Flashbots Builder.
4.2 Relationship between searchers and builders
To better interpret these results, we ask whether exclusive transactions are the outcome of exclusive relationships between senders and builders, or, instead, the same sender submits different exclusive transactions to different builders. In the latter case, the natural question is how these transactions differ across builders.
# of builders Transactions Fees (ETH)
From To From $\geq$5 To $\geq$5 From To From $\geq$5 To $\geq$5
1
3,338
(63.85%)
3,053
(58.4%)
2,757
(60.58%)
2,764
(56.35%)
80.6
(59.2%)
66.7
(49.0%)
60.55
(53.1%)
58.8
(51.6%)
2
1,356
(25.9%)
1,455
(27.83%)
1,264
(27.8%)
1,421
(28.97%)
35.1
(25.8%)
39.6
(29.1%)
32.9
(28.9%)
38.1
(33.46%)
3
281
(5.3%)
87
(1.66%)
277
(6.09%)
87
(1.77%)
15.8
(11.6%)
20.3
(14.9%)
15.8
(13.86%)
20.3
(17.8%)
4
87
(1.6%)
92
(1.76%)
87
(1.91%)
92
(1.88%)
0.85
(0.63%)
1.02
(0.75%)
0.85
(0.75%)
1.02
(0.89%)
5
166
(3.2%)
524
(10.2%)
166
(3.65%)
524
(10.68%)
3.85
(2.8%)
7.31
(5.4%)
3.85
(3.37%)
7.3
(6.41%)
8
0
(0%)
17
(0.33%)
0
(0%)
17
(0.35%)
0
(0%)
1.24
(0.91%)
0
(0%)
1.24
(1.1%)
Table 1: Exclusive transactions by sender, grouped by how many distinct builders each sender interacts with. Columns distinguish senders defined by from vs. to addresses, with and without a minimum threshold ($\geq 5$ transactions). The left panel reports transaction counts with the share of total transactions underneath. The right panel reports total fees in ETH with the share of total fees underneath.
# of builders Transactions Fees (ETH)
From To From $\geq$5 To $\geq$5 From To From $\geq$5 To $\geq$5
1
366
(74.7%)
331
(67.6%)
46
(30.1%)
162
(51.3%)
5.96
(48.6%)
0.94
(7.7%)
2.03
(24.64%)
0.87
(7.12%)
2
97
(19.8%)
123
(25.1%)
81
(52.9%)
118
(37.3%)
3.51
(28.6%)
8.47
(69.1%)
3.41
(41.37%)
8.46
(69.47%)
3
10
(2.4%)
6
(1.22%)
9
(5.9%)
6
(1.9%)
1.6
(12.9%)
1.51
(12.3%)
1.59
(19.27%)
1.51
(12.36%)
4
11
(2.24%)
11
(2.24%)
11
(7.2%)
11
(3.5%)
0.07
(0.57%)
0.07
(0.57%)
0.07
(0.85%)
0.07
(0.57%)
5
6
(1.22%)
19
(3.9%)
6
(3.9%)
19
(6.0%)
1.14
(9.3%)
1.28
(10.4%)
1.14
(13.87%)
1.28
(10.47%)
Table 2: Exclusive transactions by sender, grouped by how many distinct builders each sender interacts with. Columns distinguish senders defined by from vs. to addresses, with and without a minimum threshold ($\geq 5$ transactions). The left panel reports the count of exclusive transactions that landed on-chain, with the share of total exclusive transactions included on-chain underneath. The right panel reports total fees in ETH among the transactions that landed on-chain, with the share of total underneath.
To study the relationships between the senders of the 5,228 exclusive transactions and builders, we must first define “sender.” One option is the from address. However, many searchers use multiple from addresses while embedding their logic in smart contracts that automatically generate transactions. For searchers, therefore, a more reliable identifier is the to address, corresponding to the address of the smart contract used by the searcher. We may also want to restrict attention to senders with a minimum number of transactions. Otherwise, a sender with only one or two exclusive transactions will mechanically appear to work exclusively with a single builder, which may reflect limited data rather than a genuine relationship.
Tables 1 and 2 group transactions by the number of builders each sender interacts with, under different definitions of “sender.” Tables 1 reports the resulting transaction counts and the total fees paid by those transactions, while Table 2 reports the same but only for transaction that are ultimately included on chain. While transaction counts vary somewhat with the definition of “sender,” the fee data in Table 1 are more robust: by value, roughly half of all exclusive transactions arise from exclusive relationships between senders and builders. However, once we restrict to on-chain transactions, this share drops sharply. Using to addresses—which better identify searchers—only around 7% of the value of exclusive transactions that land on-chain can be traced to exclusive relationships. Taken together, these results suggest that exclusive relationships between senders and builders play only a limited role in explaining why, as noted earlier, 85% of the value of on-chain transactions comes from exclusives.
Table 3 performs a similar analysis as Table 1, but builder by builder. It reports the number of exclusive transactions each builder receives and the extent to which these come from senders who interact only with that builder (under different definitions of “sender”). The key insight is that the importance of exclusive relationships varies, both in absolute numbers and in the percentage of exclusive transactions received: the number of exclusive relationships goes from at least 15 (Titan builder) to zero; the fraction of exclusive transactions from exclusive relationships goes from 100% (Ty for the Block) to, again, zero.
Builder All senders Senders with $\geq$5 exclusive txs
Exclusive
txs
From
(adds/txs)
To
(adds/txs)
Exclusive
txs
From
(adds/txs)
To
(adds/txs)
Titan Builder 3106 360 / 2519 181 / 2457 2891 20 / 2115 15 / 2242
Rsync Builder 1335 79 / 220 20 / 50 1305 13 / 105 1 / 14
Ty for the Block 304 3 / 304 5 / 304 300 1 / 300 3 / 300
Boba Builder 148 9 / 146 7 / 144 137 2 / 136 1 / 136
Builder0x69 69 9 / 15 5 / 7 59 1 / 5 0 / 0
Flashbots 61 17 / 19 10 / 10 48 0 / 0 0 / 0
Gambit Labs 54 3 / 51 3 / 51 52 1 / 49 1 / 49
Bob the Builder 31 5 / 11 2 / 7 26 1 / 6 1 / 6
I Can Haz Block? 31 1 / 24 0 / 0 31 1 / 24 0 / 0
Blockbeelder.com 30 0 / 0 0 / 0 30 0 / 0 0 / 0
Geth Imposter 21 3 / 21 3 / 21 17 2 / 17 2 / 17
Beaverbuild.org 14 0 / 0 0 / 0 13 0 / 0 0 / 0
Buildernet
11 2 / 2 0 / 0 10 0 / 0 0 / 0
Payload.de 7 3 / 3 0 / 0 4 0 / 0 0 / 0
Smithbot.xyz 2 2 / 2 2 / 2
Mars 2 0 / 0 0 / 0 2 0 / 0 0 / 0
Bloxroute 1 0 / 0 0 / 0 1 0 / 0 0 / 0
Penguinbuild.org 1 1 / 1 0 / 0 1 0 / 0 0 / 0
Table 3: Number of exclusive transactions by builder, along with the number of distinct from and to addresses that sent exclusive transactions only to that builder and the number of transactions they sent. The left panel includes all senders; the right panel restricts to senders with at least 5 exclusive transactions.
We next examine how exclusive transactions from the same sender differ when sent to multiple builders. Specifically, we identify transactions originating from the same to address but submitted to different builders within the same auction cycle. [^12] As noted, multiple from addresses may use the same to address. Defining a sender by its to address, therefore, makes it more likely that we detect similarities among transactions from the same sender but with different hashes. In this sense, using to addresses provides a stricter test of our hash-based classification of transactions as exclusive, private, or public. We then re-execute these transactions to extract their logs—the sequence of contract calls and outputs each transaction generates. In total, we identify 665 such transactions producing 279 unique logs. Nearly all of them (630 transactions) are swaps in DEX pools, primarily on Uniswap V2, Uniswap V3, Sushiswap, Pancakeswap, or Curve.
We categorize these transactions based on their logs and fees:
  • Partially exclusive: An exclusive transaction is deemed partially exclusive if another exclusive transaction from the same sender in the same auction cycle produces an identical log. To measure its “degree of exclusivity,” we compare all such transactions (same sender, same cycle, same log) and compute the difference between the highest and second-highest fee paid to different builders. Note that the degree of exclusivity can be zero if multiple exclusive transactions share the same log and fee. [^13] Two transactions may generate different hashes despite being functionally identical. In our dataset, we find cases of transactions with identical logs and fees but differing in gas limit (55 cases), max priority fee (11 cases), nonce or sender (29 cases), and transaction value (32 cases, “transaction value” is an additional ETH transfer that a transaction may perform).
  • Similar: Two exclusive transactions are classified as similar if they come from the same sender in the same auction cycle, swap in the same DEX pools, but differ in the amounts swapped, resulting in distinct logs.
We find 171 partially exclusive transactions (59 unique logs) with a degree of exclusivity equal to zero—that is, the sender submitted the same transaction (same log, same fee) to at least two builders. We found another 494 partially exclusive transactions (220 unique logs) with different fees across builders. Among these, the average degree of exclusivity is 0.004178 ETH (or 23% of the value of the second-highest fee).
Finally, we identify 1,179 “similar” exclusive transactions grouped into 162 log sets (861 unique logs), where transactions differ only in the swap amount. Comparing the fees attached to these transactions, we find that the average gap between the highest payment to one builder and the second-highest payment to another builder is almost 70% of the second-highest payment. This suggests that the differences between these transactions are economically meaningful.
The takeaway is that not all transactions that appear “exclusive” at the hash level remain exclusive once we examine their logs: some are duplicates submitted to multiple builders, while others are variations of the same strategy. Quantitatively, however, these cases are limited—only 665 partially exclusive transactions out of 5,576 in total. Also, if we change the classification of partially exclusive and similar transactions to “private”, we find that 77.2% of fees paid by transactions that are included on-chain are from exclusive transactions (vs 84% with our initial classification). The main result, therefore, stands: exclusive transactions remain the key driver of builders’ profitability.
5 Conclusion
This paper provides the first systematic analysis of non-winning Ethereum blocks, allowing us to distinguish between public, private, and exclusive transactions. We show that, although exclusive transactions represent a minority of all transactions, they account for the vast majority of builder revenues—roughly 84% of fees in winning blocks. Exclusive relationships between senders and builders are widespread but account for a minimal share (approximately 7%) of the value of exclusive transactions included on-chain. These findings help explain why block-building is so concentrated, with a small number of builders dominating the market.
We also find that a significant share of transactions that appear exclusive in one bidding cycle become private or even public in later cycles. Examining transaction logs reveals that some “exclusive” transactions are actually duplicates or variations of the same strategy sent to multiple builders. Yet these cases are quantitatively limited: exclusivity measured at the hash level still provides an accurate picture of the determinants of builders’ profitability.
Our results contribute to the debate on the extreme concentration of Ethereum’s builder market. By showing that exclusive transactions generate the vast majority of builder revenues, we provide support to the view that exclusivity is an important driver of concentration and helps explain the dominance of a few builders.